# From Theory to Reality: A Design Framework for Integrated Communication and Computing Receivers

### v1

@@ I Introduction
To this end, we first formulate a system model in which the communication and computing signals are transmitted simultaneously by the user equipments (UEs) or edge devices (EDs), both of which are to be detected by the receiving base station (BS) or access point (AP). We then derive the relevant message passing rules to extract the separate data symbol and computing streams via the GaBP framework, which is initialized with a closed-form solution to an optimal combiner design problem with successive interference cancellation (SIC).
@@ III Proposed Integrated Communications and Computing Receiver Design Framework
Since the goal of AirComp is to estimate a certain function that depends on the computing signal $\bm{s}$, the assumption that the mean of the elements $\mu_{s}$ is known is restrictive, motivating the development of a closed form initialization which is elaborated on in the latter sections.
@@ III-A3 Soft Replica Generation
Similarly, since the computing signal follows a Gaussian distribution, the denoiser with a Gaussian prior and its corresponding variance is given as  [^3] It is worth noting that if the post-processed computing signals have zero mean, the Gaussian denoiser will no longer require a closed-form initialization; $i.e.$ the last three lines under the initialization section in Algorithm 1 can be ignored.
@@
$\hat{d}_{k}=\left(\sum_{n=1}^{N}\frac{|h_{n,k}|^{2}}{\tilde{\sigma}_{d:n,k}^{2(i)}}\right)^{\!\!\!-1}\!\!\left(\sum_{n=1}^{N}\frac{h^{*}_{n,k}\cdot\tilde{y}_{d:n,k}^{(i)}}{\tilde{\sigma}_{d:n,k}^{2(i)}}\right),$ (17a)
$\hat{s}_{k}=\Re\Bigg{\{}\left(\sum_{n=1}^{N}\frac{|h_{n,k}|^{2}}{\tilde{\sigma}_{s:n,k}^{2(i)}}\right)^{\!\!\!-1}\!\!\left(\sum_{n=1}^{N}\frac{h^{*}_{n,k}\cdot\tilde{y}_{s:n,k}^{(i)}}{\tilde{\sigma}_{s:n,k}^{2(i)}}\right)\Bigg{\}},$ (17b)
@@
III-B Closed-form Initialization
For the computation of a target function $\sum s_{k}$ at the BS/AP, let us first reformulate the combining of the residual signal at the $i$-th iteration of the algorithm leveraging equation (4) after SIC of the estimated communication signal $\hat{\bm{d}}^{(i)}$ as
$\tilde{f}^{(i)}(\bm{s})=\bm{u}^{\mathsf{H}}(\bm{y}-\bm{H}\hat{\bm{d}}^{(i)})=\bm{u}^{\mathsf{H}}(\bm{H}(\bm{s}-\check{\bm{d}}^{(i)})+\bm{w}),$ (18)
@@ III-B Closed-form Initialization
$\vspace{-1ex}\underset{\bm{u}\in\mathbb{C}^{N\times 1}}{\mathrm{minimize}}\hskip 12.91663pt\|f(\bm{s})-\tilde{f}^{(i)}(\bm{s})\|_{2}^{2},$ (19)
@@
$\vspace{-1ex}\|f(\bm{s})\!-\!\tilde{f}^{(i)}(\bm{s})\|_{2}^{2}\!\triangleq\!\|\mathbf{1}_{K}^{\mathsf{T}}\cdot\bm{s}\!-\!\bm{u}^{\mathsf{H}}(\bm{H}(\bm{s}\!-\!\check{\bm{d}}^{(i)})+\bm{w})\|_{2}^{2}.$ (20)
Then, the closed-form solution at a given $i$-th iteration for the combining vector can be derived as
$\vspace{-1ex}\bm{u}^{(i)}=(\bm{H}(\sigma_{s}^{2}\mathbf{I}_{K}+\bm{\Omega}^{(i)})\bm{H}^{\mathsf{H}}+\sigma^{2}_{w}\mathbf{I}_{N})^{-1}\cdot\bm{H}(\sigma_{s}^{2}\mathbf{I}_{K})\mathbf{1}_{K},$ (21)
where $\bm{\Omega}^{(i)}\triangleq\mathbb{E}[\check{\bm{d}}\check{\bm{d}}^{\mathsf{H}}]$ is the instantaneous variance of the data signal computed via the GaBP.
@@ III-C Joint Integrated Communication and Computing Design
We now combine the low complexity GaBP with the AirComp initialization to estimate both the data signal $\hat{\bm{d}}$ and computing function $\hat{f}(\bm{s})$. The complete pseudocode for the procedure is summarized in Algorithm 1.
@@
8: - Initialize $\bm{u}^{(0)}$ from equation (21).
9: - Compute $\tilde{f}^{(i)}(\bm{s})$ from equation (18).
10: - Set $\mu_{s}^{(0)}=\tilde{f}^{(i)}(\bm{s})/K$.
11:
12:for $i=1$ to $i_{\text{max}}$ do
13:Communication and Computing Update: $\forall n,k$
14:Compute soft IC data signal $\tilde{y}_{d:{n,k}}^{(i)}$ and its corresponding variance $\tilde{\sigma}^{2(i)}_{d:{n,k}}$ from equations (7a) and (9a).
15:Compute soft IC computing signal $\tilde{y}_{s:{n,k}}^{(i)}$ and its corresponding variance $\tilde{\sigma}^{2(i)}_{s:{n,k}}$ from equations (7b) and (9b).
16:Compute extrinsic data signal belief $\bar{d}_{n,k}^{(i)}$ and its corresponding variance $\bar{\sigma}_{d:{n,k}}^{2(i)}$ from equations (11a) and (12a).
17:Compute extrinsic computing signal belief $\bar{s}_{n,k}^{(i)}$ and its corresponding variance $\bar{\sigma}_{s:{n,k}}^{2(i)}$ from eqs. (11b) and (12b).
18:Compute denoised and damped data signal estimate $\hat{d}_{n,k}^{(i)}$ from equations (13) and (15a).
19:Compute denoised and damped data signal variance $\hat{\sigma}_{d:{n,k}}^{2(i)}$ from equations (6a) and (16a).
20:Compute denoised and damped computing signal estimate $\hat{s}_{n,k}^{(i)}$ from equations (14a) and (15b).
21:Compute denoised and damped computing signal variance $\hat{\sigma}_{s:{n,k}}^{2(i)}$ from equations (14b) and (16b).
22:Optional High Performance Update:
23:Update $\bm{u}^{(i)}$ from equation (21).
24:Update $\tilde{f}^{(i)}(\bm{s})$ from equation (18).
25:Update $\mu_{s}^{(i)}=\tilde{f}^{(i)}(\bm{s})/K$.
26:end for
27:Calculate $\hat{d}_{k},\forall k$ using equation (17a).
28:Calculate $\hat{s}_{k},\forall k$ using equation (17b).
29:Compute $\hat{f}(\bm{s})=\sum_{k=1}^{K}\hat{s}_{k}$.
@@ IV Performance Assessment
For all the numerical simulations carried out, the total transmit power is held constant at $1$ with $99\%$ of the power allocated to the data signal and $1\%$ of the power allocated to the computing signal. In addition, the computing signal is assumed to follow $s_{k}\sim\mathcal{N}(0,\sigma_{s}^{2})$ and the channel coefficients follow $h_{n,k}\sim\mathcal{CN}(0,1)$. Similarly, the algorithmic parameters are set as $\beta_{d}=0.5$, $\beta_{s}=0.8$ and $i_{\text{max}}=30$.
@@ IV-A Numerical Results
Consequently, Fig. 2 now showcases BER and MSE results for the estimation of $\hat{\bm{d}}$ and $\hat{f}(\bm{s})$, respectively, for all three aforementioned cases as follows: the linear bound is obtained by running the proposed algorithm including the high performance update under the assumption that one of the estimates is perfectly known ($i.e.,$ for the case of data detection, the computing signals are assumed to be known and vice-versa); the high performance result is computed by running the proposed algorithm with the high performance update included; the low complexity result is obtained by skipping the high performance update ($i.e.,$ lines $9-11$ in algorithm 1 are not executed) and finally, as a consequence of $s_{k}\sim\mathcal{N}(0,\sigma_{s}^{2})$, the red dashed line presents the results when the closed-form initialization is also ignored.
@@
Next, we consider a similar system setup but with a much larger number of $N=200$ antennas servicing a varying number of single-antenna users that are also categorized into the underloaded, fully loaded and overloaded scenarios with $K=50$, $K=100$ and $K=200$ users, respectively, to demonstrate the efficacy of the low complexity variation of the proposed algorithm for enabling ICC in massive scenarios.
@@ IV-B Complexity Analysis
Since the high performance update requires a matrix inversion of size $N\times N$ at every iteration, the resulting per-iteration computational complexity order is $\mathcal{O}(N^{3})$. However, since the low complexity variation only requires this matrix inversion once at the initialization stage, the resulting per-iteration computational complexity order is $\mathcal{O}(NK)$ which is linear on the number of element-wise computations performed.

### v2

@@ I Introduction
To this end, we first formulate a system model in which the communication and computing signals are transmitted simultaneously by the user equipments (UEs) or edge devices (EDs), both of which are to be detected by the receiving base station (BS) or access point (AP). We then derive the relevant message passing rules to extract the separate data symbol and computing streams via the GaBP framework, enhanced with a closed-form combiner design to an optimization problem incorporating successive interference cancellation (SIC).
@@ III Proposed Integrated Communications and Computing Receiver Design Framework
Since the goal of AirComp is to estimate a certain function that depends on the computing signal $\bm{s}$, the assumption that the mean of the elements $\mu_{s}$ is known is restrictive and has to updated iteratively with its corresponding estimates as detailed in the subsequent subsections.
@@ III-A3 Soft Replica Generation
Similarly, since the computing signal follows a Gaussian distribution, the denoiser with a Gaussian prior and its corresponding variance is given as  [^3] It is worth noting that if the post-processed computing signals have zero mean, the Gaussian denoiser will only depend on the extrinsic means and the computing symbol variance.
@@
$\hat{d}_{k}=\left(\sum_{n=1}^{N}\frac{|h_{n,k}|^{2}}{\tilde{\sigma}_{d:n,k}^{2(i_{\text{max}})}}\right)^{\!\!\!-1}\!\!\left(\sum_{n=1}^{N}\frac{h^{*}_{n,k}\cdot\tilde{y}_{d:n,k}^{(i_{\text{max}})}}{\tilde{\sigma}_{d:n,k}^{2(i_{\text{max}})}}\right),$ (17a)
$\hat{s}_{k}^{(i)}=\Re\Bigg{\{}\left(\sum_{n=1}^{N}\frac{|h_{n,k}|^{2}}{\tilde{\sigma}_{s:n,k}^{2(i)}}\right)^{\!\!\!-1}\!\!\left(\sum_{n=1}^{N}\frac{h^{*}_{n,k}\cdot\tilde{y}_{s:n,k}^{(i)}}{\tilde{\sigma}_{s:n,k}^{2(i)}}\right)\Bigg{\}},$ (17b)
@@
III-B Closed-form AirComp Combiner Design
For the computation of a target function $\sum s_{k}$ at the BS/AP, let us first reformulate the combining of the residual signal leveraging equation (4) after SIC of the estimated communication signal $\hat{\bm{d}}$ as
$\vspace{-1ex}\hat{f}(\bm{s})=\bm{u}^{\mathsf{H}}(\bm{y}-\bm{H}\hat{\bm{d}})=\bm{u}^{\mathsf{H}}(\bm{H}(\bm{s}-\check{\bm{d}})+\bm{w}),$ (18)
@@ III-B Closed-form AirComp Combiner Design
$\underset{\bm{u}\in\mathbb{C}^{N\times 1}}{\mathrm{minimize}}\hskip 12.91663pt\|f(\bm{s})-\hat{f}(\bm{s})\|_{2}^{2},$ (19)
@@
$\|f(\bm{s})\!-\!\hat{f}(\bm{s})\|_{2}^{2}\!\triangleq\!\|\mathbf{1}_{K}^{\mathsf{T}}\cdot\bm{s}\!-\!\bm{u}^{\mathsf{H}}(\bm{H}(\bm{s}\!-\!\check{\bm{d}})+\bm{w})\|_{2}^{2}.$ (20)
Then, the closed-form solution for the combining vector can be derived as
$\bm{u}=(\bm{H}(\sigma_{s}^{2}\mathbf{I}_{K}+\bm{\Omega})\bm{H}^{\mathsf{H}}+\sigma^{2}_{w}\mathbf{I}_{N})^{-1}\cdot\bm{H}\cdot(\sigma_{s}^{2}\mathbf{I}_{K})\cdot\mathbf{1}_{K},$ (21)
where $\bm{\Omega}\triangleq\mathbb{E}[\check{\bm{d}}\check{\bm{d}}^{\mathsf{H}}]$ is the final instantaneous variance of the data signal computed via the GaBP.
@@ III-C Joint Integrated Communication and Computing Design
We now combine the low complexity GaBP with the closed-form AirComp combiner to estimate both the data signal $\hat{\bm{d}}$ and obtain the computing function $\hat{f}(\bm{s})$. The complete pseudocode for the procedure is summarized in Algorithm 1.
@@
8: - Set $\mu_{s}^{(0)}=0$.
9:
10:for $i=1$ to $i_{\text{max}}$ do
11:Communication and Computing Update: $\forall n,k$
12:Compute soft IC data signal $\tilde{y}_{d:{n,k}}^{(i)}$ and its corresponding variance $\tilde{\sigma}^{2(i)}_{d:{n,k}}$ from equations (7a) and (9a).
13:Compute soft IC computing signal $\tilde{y}_{s:{n,k}}^{(i)}$ and its corresponding variance $\tilde{\sigma}^{2(i)}_{s:{n,k}}$ from equations (7b) and (9b).
14:Compute extrinsic data signal belief $\bar{d}_{n,k}^{(i)}$ and its corresponding variance $\bar{\sigma}_{d:{n,k}}^{2(i)}$ from equations (11a) and (12a).
15:Compute extrinsic computing signal belief $\bar{s}_{n,k}^{(i)}$ and its corresponding variance $\bar{\sigma}_{s:{n,k}}^{2(i)}$ from eqs. (11b) and (12b).
16:Compute denoised and damped data signal estimate $\hat{d}_{n,k}^{(i)}$ from equations (13) and (15a).
17:Compute denoised and damped data signal variance $\hat{\sigma}_{d:{n,k}}^{2(i)}$ from equations (6a) and (16a).
18:Compute denoised and damped computing signal estimate $\hat{s}_{n,k}^{(i)}$ from equations (14a) and (15b).
19:Compute denoised and damped computing signal variance $\hat{\sigma}_{s:{n,k}}^{2(i)}$ from equations (14b) and (16b).
20:Compute $\hat{s}_{k}^{(i)},\forall k$ using equation (17b).
21:Update $\mu_{s}^{(i)}=\frac{1}{K}\sum_{k=1}^{K}\hat{s}_{k}^{(i)}$.
22:end for
23:Communication and Computing Consensus:
24:Calculate $\hat{d}_{k},\forall k$ (equivalently $\hat{\bm{d}}$) using equation (17a).
25:Compute $\bm{u}$ from equation (21).
26:Compute $\hat{f}(\bm{s})$ from equation (18).
@@ IV Performance Assessment
For all the numerical simulations carried out, the total transmit power is held constant at $1$ with $99\%$ of the power allocated to the data signal and $1\%$ of the power allocated to the computing signal to minimize the distortion to the overall transmit waveform, allowing for the exploitation of all the degrees-of-freedoms (DoFs) in waveform design for communications and/or sensing. In addition, the computing signal is assumed to follow $s_{k}\sim\mathcal{N}(0,\sigma_{s}^{2})$ and the channel coefficients follow $h_{n,k}\sim\mathcal{CN}(0,1)$. Similarly, the algorithmic parameters are set as $\beta_{d}=0.5$, $\beta_{s}=0.8$ and $i_{\text{max}}=30$.
@@ IV-A Numerical Results
Consequently, Fig. 2 now showcases BER and MSE results for the estimation of $\hat{\bm{d}}$ and $\hat{f}(\bm{s})$, respectively, for all three aforementioned cases as follows: the linear bound is obtained by running the proposed algorithm under the assumption that one of the estimates is perfectly known ($i.e.,$ for the case of data detection, the computing signals are assumed to be known and vice-versa); the result illustrated by the white circles is computed by running the proposed algorithm as it is and finally, the red dashed line presents the results under the assumption that the computing signals follow $s_{k}\sim\mathcal{N}(0,\sigma_{s}^{2})$, leading to $\mu_{s}=0$.
@@
Next, we consider a similar system setup but with a much larger number of $N=200$ antennas servicing a varying number of single-antenna users that are also categorized into the underloaded, fully loaded and overloaded scenarios with $K=50$, $K=100$ and $K=200$ users, respectively, to demonstrate the efficacy of the proposed algorithm for enabling ICC in massive scenarios.
@@ IV-B Complexity Analysis
Due to the low complexity nature of the GaBP framework, the resulting per-iteration computational complexity order for the initial communication and computing update is $\mathcal{O}(NK)$ which is linear on the number of element-wise computations performed. However, since the final closed-form computing combiner on line 12 of Algorithm 1 requires a matrix inversion of size $N\times N$, the computational complexity order for the latter part is $\mathcal{O}(N^{3})$.